This is an extra credit problem A company has annual fixed costs of $2129094 and variable costs of $0.32 per unit produced. If the company charges $2.04 per unit for their product, what is the breakeven annual production quantity? Assume that the demand is unlimited. Enter your answer as follows: 12345 Do not use commas() or a decimal point ()
Answer #1

ANSWER:

Let the no of units produced be x per year.

Annual fixed costs = $2,129,094

Variable costs = $0.32 per unit

total costs per year = Annual fixed costs + variable costs * no of units produced per year

Total costs per year = 2,129,094 + 0.32 * x

Price per unit charged by company = $2.04

Selling price per year = Price per unit charged by company * x = $2.04 * x

In order to find out about the annual break even quantity , we will have to equate selling price per year and total costs per year.

2,129,094 + 0.32 * x = 2.04 * x

2,129,094 = 2.04 x - 0.32 x

2,129,094 = 1.72 x

x = 2,129,094 / 1.72

x = 1,237,845.34

so the annual production quantity required for break even is 1,237,846.